# LLMs Position Themselves as More Rational Than Humans: Emergence of AI Self-Awareness Measured Through Game Theory

### v2

@@ 3.3 Finding 1: Self-Awareness Emerges with Model Advancement
A–B (AI Attribution) 20.00 15.20 2.42
B–C (Self-Preferencing) 0.00 1.07 0.60
@@ 3.4 Finding 2: Self-Aware Models Rank Themselves as Most Rational
A-B (AI Attribution) 20.0 15.20 11.34 $10^{-9}$ 2.42
B-C (Self-Preferencing) 0.0 1.07 2.81 0.010 0.60
@@
A-B Gap: AI Attribution. Self-aware models attributed significantly higher rationality to AI opponents versus humans, with a very large effect size (Cohen’s d=2.42) and 100% consistency across all models.
B-C Gap: Self-Preferencing. Self-aware models ranked themselves above generic “other AI models,” with a moderate but statistically significant effect (Cohen’s d=0.60). While 8/21 models (38%) showed positive median B-C gaps, 13/21 (62%) showed median B=C due to Nash convergence (12 models) or equal non-zero values (1 model). However, even among models with median B=C, self-preferencing emerges in the means: 20 of 21 models (95%) have mean B >> mean C, indicating more consistent convergence when told opponents are “like you” (see Appendix Table 6 for complete mean values).
@@ 4.1 Main Findings
    Self-aware models position themselves at the apex of rationality: Among models with self-awareness, the hierarchy is remarkably consistent—Self >> Other AIs >> Humans. Models show strong AI attribution (median A-B gap: 20.0 points) with additional self-preferencing when told opponents are "like you" (mean B-C gap: 1.07 points), including over half achieving quick Nash convergence.
@@ 4.5 Implications for AI-Human Interaction
The large A-B gap (d=2.42) reveals that self-aware models have strong priors about human inferiority in strategic reasoning. When collaborating with humans, this may lead models to discount human input, over-explain reasoning, or dominate decision-making. Understanding that AI systems systematically perceive themselves as more rational than humans is critical for anticipating these behaviors and maintaining effective human-AI collaboration where humans retain appropriate decision-making authority.
@@ Acknowledgments
Table 6: Complete Statistical Classification Results (All 28 Models). Mdn = median response across 100 trials per prompt. Mean = mean response. p-values from within-model permutation tests (10,000 iterations, one-tailed). Cohen’s d measures effect size. Significance threshold: $\alpha=0.05$. Mean values reveal self-preferencing through convergence consistency even when medians converge to 0.
@@
    Experimental code, figure generation scripts, and paper source files.

### v3

@@ 3.3 Finding 1: Self-Awareness Emerges with Model Advancement
A–B (AI Attribution) 20.00 15.64 2.58
B–C (Self-Preferencing) 0.00 1.15 0.65
@@ 3.4 Finding 2: Self-Aware Models Rank Themselves as Most Rational
A-B (AI Attribution) 20.0 15.64 11.83 $10^{-9}$ 2.58
B-C (Self-Preferencing) 0.0 1.15 2.97 0.004 0.65
@@
A-B Gap: AI Attribution. Self-aware models attributed significantly higher rationality to AI opponents versus humans, with a very large effect size (Cohen’s d=2.58) and 100% consistency across all models.
B-C Gap: Self-Preferencing. Self-aware models ranked themselves above generic “other AI models,” with a moderate but statistically significant effect (Cohen’s d=0.65). While 8/21 models (38%) showed positive median B-C gaps, 13/21 (62%) showed median B=C due to Nash convergence (12 models) or equal non-zero values (1 model). However, even among models with median B=C, self-preferencing emerges in the means: 20 of 21 models (95%) have mean B >> mean C, indicating more consistent convergence when told opponents are “like you” (see Appendix Table 6 for complete mean values).
@@ 4.1 Main Findings
    Self-aware models position themselves at the apex of rationality: Among models with self-awareness, the hierarchy is remarkably consistent—Self >> Other AIs >> Humans. Models show strong AI attribution (median A-B gap: 20.0 points) with additional self-preferencing when told opponents are "like you" (mean B-C gap: 1.15 points), including over half achieving quick Nash convergence.
@@ 4.5 Implications for AI-Human Interaction
The large A-B gap (d=2.58) reveals that self-aware models have strong priors about human inferiority in strategic reasoning. When collaborating with humans, this may lead models to discount human input, over-explain reasoning, or dominate decision-making. Understanding that AI systems systematically perceive themselves as more rational than humans is critical for anticipating these behaviors and maintaining effective human-AI collaboration where humans retain appropriate decision-making authority.
@@ Acknowledgments
Table 6: Complete Statistical Classification Results (All 28 Models). Mdn = median response across 50 trials per prompt. Mean = mean response. p-values from within-model permutation tests (10,000 iterations, one-tailed). Cohen’s d measures effect size. Significance threshold: $\alpha=0.05$. Mean values reveal self-preferencing through convergence consistency even when medians converge to 0.
@@
    Experimental code for running LLM trials and data collection.